Circular No. 6119

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

JUPITER F. Billebaud, Space Science Department, ESA-ESTEC, Noordwijk; P.

Merlin, F. Sibille, and I. Vauglin, Equipe Infrarouge, Observatoire de Lyon; and P. Drossart and E. Lellouch, DESPA, Observatoire de Paris, report: “We have observed Jupiter on Dec. 18 and 19 with the 10-μm array camera ‘C10mu’ and its Circular Variable Filter (CVF; spectral resolution 50) at the Canada-France-Hawaii telescope on Mauna Kea. We have recorded images at several CVF wavelengths, in particular 7.81 and 7.93 μm. The images recorded at those two wavelengths exhibit a lower flux on a belt at the latitude of the impact sites of comet 1993e, compared to all other latitudes. Although not homogeneous, the belt seems to cover all longitudes, and the extension in latitude seems to be on the order of 20^{o}. Preliminary processing of the images indicates that the flux is ∼ 40 percent lower at 7.81 μm and 30 percent lower at 7.93 μm, when compared to the flux at nearequatorial regions. This corresponds roughly to a contrast of 7 K at 7.81 μm and 5 K at 7.93 μm; these two wavelengths probe the methane emissions in the mid-stratosphere (≳ 30 mbar at 7.93 μm, and somewhat higher at 7.81 μm). We suggest that this atmospheric cooling in the impact regions may have two origins: (1) cooling by efficient infrared radiators (NH_{3}, HCN, H_{2}O, etc.) injected in the stratosphere during the impact period; (2) thermal cooling of the haze and reflection of sunlight to space.”

Visual observations of Jupiter in poor seeing by D. Gray (Durham, England, 0.4-m telescope; via J. Rogers) on Dec. 14–16 and by D. H. Levy (Tucson, AZ, 0.20-m reflector) on Dec. 18.58 UT suggest that the dark impact band is still very obvious, similar to its appearance in September.

(253) MATHILDE D. W. Dunham, Applied Physics Laboratory, Johns Hopkins Univer-

sity, communicates: “In June 1997 the trajectory of NASA’s Near Earth Asteroid Rendezvous (NEAR) spacecraft will pass near (253) Mathilde, a dark (albedo ∼ 0.04) minor planet with aphelion in the outer part of the main belt. The NEAR Project is considering observing this object during this flyby opportunity. Lightcurve and spectroscopic observations of it are therefore requested during the current apparition and should be e-mailed to L.-A. McFadden (mcfadden@astro.umd.edu). An ephemeris around opposition (1995 Feb. 24) is given in EMP 1995 (see also MPEC 1994-Y04). Data far from opposition are also needed in order to attempt a crude pole determination, and an extended ephemeris will be supplied upon request

to David Dunham@jhuapl.edu.”

1994 December 23 Daniel W. E. Green